The purchase price for a company car is $19,800. The useful life is 6 years after which it will have an estimated scrap value of $1,200.
-Find the linear equation relating book value and number of years.
- What is the book value after four years?
Show all work.

Answer by Theo(13342)

the price of the car is 19,800.
that's the book value at the end of year 0.
the end of year 0 is the beginning of year 1.
the salvage value at the end of year 6 is 1,200.
that's the book value at the end of year 6.

using straight line depreciation, the formula would be:
y = mx + b
m is the slope
b is the y-intercept
y is the value after x years.
x is the number of years.

when x = 0, y = 19,800.
when x = 6, y = 1,200

to find the slope, you need 2 points on the line.
the first point is (x1,y1) = (0,19800)
the second point is (x2,y2) = (6,1200)

the slope is (y2-y1)/(x2-x1) = (1200 - 19800)/(6-0) = -18600/6 = -3100
the equation becomes y = -3100 * x + b
replace x and y with one of the points to find b.
i used (0,19800)
y = -3100 * x + b becomes 19800 = -3100 * 0 + b
solve for b to get:
b = 19800.
the equation becomes y = -3100 * x + 19800

when x = 0, the equation becomes y = -3100 * 0 + 19800.
solve for y to get y = 19800.
that's the book value at the end of year 0.

when x = 6, the equation becomes y = -3100 * 6 + 19800.
solve for y to get y = 1200.
that's the book value at the end of year 6.

when x = 4, the equation becomes y = -3100 * 4 + 19800.
solve for y to get y = 7400.
that's the book value at the end of year 4.
that's your solution.